A Carnot air conditioner operates between an indoo
Answer #1

W = Q(T1/T2 - 1) { T1 = 273 + 40 , T 2 = 20 +273 , Q is heat transfer and W is amount of hea required}

W = 2000( 312/293 -1 )

W = 130 J